The length of a rectangle is 2cm longer than it's width. If the perimeter of the rectangle is 44 cm, find it's area
Answer by JulietG(1812)

Perimeter is a fence around. P = 2L + 2W
L = W + 2
Take the known value of the length from the second equation and substitute it into the first.
P = 2L + 2W
44 = 2L + 2W
44 = 2(W+2) + 2W
Distribute
44 = 2W + 4 + 2W
44 = 4W + 4
Subtract 4 from each side
40 = 4W
Divide each side by 4
10 = W
.
If the width is 10, the length is 10+2, or 12
10 + 10 (two widths) + 12 + 12 (two lengths) = 44
Success!